Calculate the work done in displacing a body with a force,f=(30î + 40j)N through S=(2tî-4tj)m between t1=2 second and t2=5 seconds


1
Expert's answer
2021-01-29T02:25:30-0500

By the definition of the work done, we have:


"W=Fs"

Let's first find displacements of the body at time "t_1=2\ s" and "t_2=5\ s":


"d_1=(2\cdot2\hat{i}-4\cdot2\hat{j})=(4\hat{i}-8\hat{j})\ m,""d_2=(2\cdot5\hat{i}-4\cdot5\hat{j})=(10\hat{i}-20\hat{j})\ m."

The displacement of the body between "t_1=2\ s" and "t_2=5\ s" can be calculated as follows:


"d=d_2-d_1=(10\hat{i}-20\hat{j})-(4\hat{i}-8\hat{j})=(6\hat{i}-12\hat{j})\ m."

Finally, we can find the work done between "t_1=2\ s" and "t_2=5\ s":


"W=(30\hat{i}+40\hat{j})\ N\cdot(6\hat{i}-12\hat{j})\ m,""W=(30, 40)\cdot(6,-12)=30\cdot6+40\cdot(-12)=-300\ J."